OpenStudy (anonymous):

A 110-g bullet is fired from a rifle having a barrel 0.636 m long. Choose the origin to be at the location where the bullet begins to move. Then the force (in newtons) exerted by the expanding gas on the bullet is 13500 + 11000x - 25750x2, where x is in meters. (a) Determine the work done by the gas on the bullet as the bullet travels the length of the barrel. ____kJ (b) If the barrel is 1.02 m long, how much work is done? ___kJ (c) How does this value compare with the work calculated in part (a)? percent difference = ___%

OpenStudy (anonymous):

\[\int\limits_{0}^{1.02} (13500 + 11000x - 25750x^2)dx\]

OpenStudy (anonymous):

= Work

OpenStudy (anonymous):

\[ \int\limits_{0}^{.63}(13500 + 11000x - 25750x^2)\]

OpenStudy (anonymous):

for both cases

OpenStudy (anonymous):

i did that its the percent difference that im kinda confused

OpenStudy (anonymous):

divide one amount by the other.... subtract 1 from the result and then multiply by 100 that's the % change eg increase from 100 to 120 .... 120/100 = 1.2; 1.2-1 =.2; .2*100 =20%

OpenStudy (anonymous):

okk i got it thank you !